# Convergence in probability of numerical solutions of a highly non-linear delayed stochastic interest rate model

### v1

@@ Abstract
We examine a delayed stochastic interest rate model with super-linearly growing coefficients and develop several new mathematical tools to establish the properties of its true and truncated EM solutions. Moreover, we show that the true solution converges to the truncated EM solutions in probability as the step size tends to zero. Further, we support the convergence result with some illustrative numerical examples and justify the convergence result for the Monte Carlo evaluation of some financial quantities.
@@ 1 Introduction
Stochastic modelling of interest rates is important for calibrating and valuing financial products such as option contracts. There are various stochastic interest rate models proposed by some authors in the existing literature to describe the evolution of interest rates. One of the well-known stochastic interest rate models, also known as the CIR model, was proposed by Cox, Ingersoll, and Ross in [5]. This model is described by the dynamics
@@
Here, $x=(x(t),t\geq 0)$ denotes the interest rate with initial value $x(0)=x_{0}$, $\alpha,\mu,\sigma>0$, and $B=(B(t),t\geq 0)$ is a scalar Brownian motion. The CIR model is mean-reverting, and due to its square root diffusion factor, it can also avoid possible negative rates.
We observe in SDE (1) that the volatility term $\sigma$ is assumed constant. However, as supported by empirical studies, volatility is not constant but exhibits empirical features widely known as volatility skews and smiles that are prevalent in option markets (see, e.g., [7, 6]). In order to capture the evolution of volatility skews and smiles, some authors proposed to model volatility as a function of a delay variable. For instance, Arriojas et al. in [1] proposed the delayed Black-Scholes model where both drift and diffusion terms are functions of a delay variable. This model is described by the dynamics
@@
on $t\in[0,T]$ with the initial value $\xi(t)\in[-\tau,0]$, where $\xi:\Omega\rightarrow C([-\tau,0]:\mathbb{R})$, $\mu,a,b>0$, $\tau=\max(a,b)$, x(ta)x(t-a) and x(tb)x(t-b) denote delays in $x(t)$, $g:\mathbb{R}\rightarrow\mathbb{R}$ is a continuous function, and $B=(B(t),t\geq 0)$ is a scalar Brownian motion. The authors showed that the delayed Black-Scholes model maintains the no-arbitrage property and the completeness of the market with the correct volatility skews and smiles. Furthermore, Mao and Sabanis in [13] introduced the delay geometric Brownian motion described by the dynamics
@@
on $t\geq 0$ with the initial value $\xi(t)$ on $t\in[-\tau,0]$, where $\tau,r>0$, $\xi:\Omega\rightarrow C([-\tau,0]:\mathbb{R})$, $B=(B(t),t\geq 0)$ is a scalar Brownian motion, the volatility function $v$ depends on $x(t-\tau)$, and $x(t-\tau)$ denotes delay in $x(t)$. The authors studied the quantitative properties of this model where $v$ is locally Lipschitz continuous and bounded. The authors provided numerical evidence to justify that the system of type (3) is a rich alternative model for an asset price process in a complete market characterised by volatility skews and smiles. The reader may also consult (e.g., [16, 9]) for financial models with features of past dependency.
In recent years, more empirical studies have shown that the most successful continuous-time models for interest rates are those models that allow the volatility of interest changes to be highly sensitive to the level of the rates (see, e.g., [3, 14]). This motivated Wu et al. in [15] to extend SDE (1) to include a super-linear diffusion term described by the dynamics
@@
on $t\geq 0$, where $\theta>1$. One notable unique feature of SDE (4) is that the solution $x(t)$ is a highly sensitive mean-reverting process. The authors established the convergence in probability of the EM solutions to the true solution. They justified the convergence result within Monte Carlo simulations to value the expected payoff of a bond and a barrier option. To further capture volatility skews and smiles and high non-linearities in the rate, the authors in [4] extended the generalised Ait-Sahalia interest rate model to include a volatility as a function of a delay variable described by the dynamics
@@
on $t\geq 0$ with the initial value $\xi(t)$ on $t\in[-\tau,0]$, where $\alpha_{-1},\alpha_{0},\alpha_{1},\alpha_{2},\tau>0$, $\xi:\Omega\rightarrow C([-\tau,0]:\mathbb{R})$, $\gamma,\theta>1$ and $B=(B(t),t\geq 0)$ is a scalar Brownian motion, $v$ is a function of $x(t-\tau)$ and $x(t-\tau)$ denotes delay in $x(t)$. Under a monotone condition and the assumption that $v$ is locally Lipschitz continuous and bounded, the authors proved the strong convergence of the truncated EM solutions to the true solution of SDDE (5) and justified that the convergence result can be applied to value a bond and a barrier option.
Therefore, in order to account for high non-linearities in the rates as well as the evolution of volatility skews and smiles, we consider it necessary to reformulate SDE (4) as SDDE with super-linearly growing drift and diffusion coefficients described by the dynamics
@@
on $t\geq 0$ with the initial value $\xi(t)$ on $t\in[-\tau,0]$, where $\tau>0$, $\xi:\Omega\rightarrow C([-\tau,0]:\mathbb{R})$, $\theta,r>0$, and $\gamma>1$. We observe that both the drift factor $x^{\gamma}$ and the diffusion factor $x^{\theta}y^{r}$ of SDDE (6) are growing super-linearly and thus violate the global Lipschitz and linear growth conditions. This is further complicated by the presence of the unbounded delay variable $y$. Therefore, it can be very challenging to obtain the solution of SDDE (6) by an analytical closed-form formula. To the best of our knowledge, there exists no relevant literature for the numerical analysis of SDDE (6) either in the strong sense or weak sense. In this case, we recognise the need to examine the feasibility of the system of SDDE (6) from a viewpoint of financial applications. Therefore, we need an efficient numerical method with fast computational performance to estimate the solution. However, in most real-world applications, the explicit EM method is preferred to the implicit type due to its simple algebraic structure, cheap computational cost, and acceptable convergence rate. It is well-known in [8] that the explicit EM scheme diverges in the strong mean-square sense at a finite point for SDEs with super-linearly growing coefficient terms. In this work, we aim to construct a variant of the truncated EM method developed in [11] to estimate the true solution of SDDE (6) and show that the truncated EM solutions converge to the true solution in probability when the step size is sufficiently small. The remainder of the paper is organised as follows: We explore mathematical notations in Section 2. In Section 3, we study properties of the true solution of SDDE (6). We construct the truncated EM techniques to approximate SDDE (6) and study properties of the truncated EM solutions in Section 4. In Section 5, we show that the truncated EM solutions converge to the true solution of SDDE (6) in probability. We also provide illustrative numerical examples to support the convergence result and justify the result via the efficient use of the Monte Carlo method to value a bond and a lookback put option in this section.
@@ 2 Mathematical preliminaries
There exist constants $D>0$ and $\ell\in(0,1]$ such that for all $\tau\leq s\leq t\leq 0$, the initial value $\xi$ satisfies
@@
We introduce the following important lemma for later use.
@@
Let Assumption 2.1 hold. For any $R>0$, there exists a positive constant $G_{R}$ such that the coefficients of SDDE (6) fulfil
$|f(x)-f(\bar{x})|+|g(x,y)-g(\bar{x},\bar{y})|\leq G_{R}(|x-\bar{x}|+|y-\bar{y}|)$ (12)
for all $x,y,\bar{x},\bar{y}\in\mathbb{R}$ with $|x|\vee|\bar{x}|\vee|y|\vee|\bar{y}|\leq R$.
@@
If we assume that $x<\bar{x}$ and $y<\bar{y}$, then
$(f(x)-f(\bar{x}))+(g(x,y)-g(\bar{x},\bar{y}))\leq(\alpha(\mu-x^{\gamma})-\alpha(\mu-\bar{x}^{\gamma}))+(\sigma x^{\theta}y^{r}-\sigma\bar{x}^{\theta}\bar{y}^{r}).$
However, we note from the Young inequality that
[…]
It then follows that
$(f(x)-f(\bar{x}))+(g(x,y)-g(\bar{x},\bar{y}))\leq-\alpha(x^{\gamma}-\bar{x}^{\gamma})+\sigma(x^{2\theta}-\bar{x}^{2\theta})+\sigma(y^{2r}-\bar{y}^{2r}).$
So by the mean-value theorem, we now get
$|f(x)-f(\bar{x})|+|g(x,y)-g(\bar{x},\bar{y})|\leq$ $-\alpha\gamma(|x^{\gamma-1}|+|\bar{x}^{\gamma-1}|)|x-\bar{x}|$
$+2\theta\sigma(|x^{2\theta-1}|+|\bar{x}^{2\theta-1}|)|x-\bar{x}|$
$+2r\sigma(|y^{2r-1}|+|\bar{y}^{2r-1}|)|y-\bar{y}|.$
Noting that
$\gamma-1>2r+2\theta-2\Rightarrow\gamma+1>2(r+\theta),$
we can apply Assumption 2.1 to obtain
$|f(x)-f(\bar{x})|+|g(x,y)-g(\bar{x},\bar{y})|\leq G_{R}(|x-\bar{x}|+|y-\bar{y}|),$
where $G_{R}\geq-\alpha\gamma(|x^{\gamma-1}|+|\bar{x}^{\gamma-1}|)+2\theta\sigma(|x^{2\theta-1}|+|\bar{x}^{2\theta-1}|))+2r\sigma(|y^{2r-1}|+|\bar{y}^{2r-1}|)$. ∎
@@ 3.1 Existence of positive solution
As $k\rightarrow\infty$, we have $\eta_{\infty}>T$ almost surely. Repeating this procedure for $t_{1}\in[0,\infty)$, it is easy to see that $\mathbb{P}(\eta_{\infty}\leq\infty)\rightarrow 0$ as $k\rightarrow\infty$. This implies that $\eta_{\infty}=\infty$ almost surely and hence we must have $\eta_{e}=\infty$ almost surely as the required assertion. ∎
@@ 3.2 Boundedness
$\mathbb{P}(/1k<x(t)<k)<1-\Big[\xi(0)^{\beta}-1-\beta\log(\xi(0))+\bar{K}_{0}\Big]\Big[\frac{1}{(1/k)^{\beta}-1+\beta\log(k)}+\frac{1}{k^{\beta}-1-\beta\log(k)}\Big],$
@@ 5 Convergence analysis
In this section, we study the finite-time convergence of the truncated EM solutions to the true solution of SDDE (6). Further, we show that the truncated EM solutions converge to the true solution of SDDE (6) in probability. We perform simulation examples to support the convergence result and justify the convergence result to evaluate some option contracts.
@@ 5.2 Convergence in probability
$\frac{2V(\xi(0))+K_{0}T+K_{1}T}{V(k)\wedge V(1/k)}<\frac{\epsilon}{2}$ (39)
@@
$\frac{K_{3}D\Delta^{\ell}+c_{p}(K_{2}+K_{3})\Delta^{1/2}\psi(\Delta)T}{V(1/k)\wedge V(k)}+\frac{K_{8}(\Delta^{\ell}\vee\Delta^{p/2}(\psi(\Delta))^{p})}{\lambda}<\frac{\epsilon}{2}.$ (40)
@@
$\mathbb{P}\Big(\sup_{0\leq t\leq T}|x_{\Delta}(t)-x(t)|^{p}\geq\lambda\Big)<\epsilon,$ (41)
@@ Acknowledgement
The author would like to acknowledge the financial support from the Heilbronn Institute for Mathematical Research (HIMR) and the UKRI/EPSRC Additional Funding Programme for Mathematical Sciences.

### v2

@@ Abstract
We examine a delayed stochastic interest rate model with super-linearly growing coefficients and develop several new mathematical tools to establish properties of its true and truncated EM solutions. Moreover, we show that the true solution converges to the truncated EM solutions in probability as the step size tends to zero. Further, we support the convergence result with some illustrative numerical examples and justify the result via efficient use of the Monte Carlo method to value some financial quantities.
@@ 1 Introduction
Stochastic modelling of interest rates is important for calibrating and valuing financial products such as option contracts. There exist variety of stochastic interest rate models proposed by some authors in the existing literature to describe the evolution of interest rates. One of the well-known stochastic interest rate models, also known as the CIR model was proposed by Cox, Ingersoll and Ross in [5]. This model follows the dynamics
@@
Here $x=(x(t),t\geq 0)$ denotes the interest rate with initial value $x(0)=x_{0}$, $\alpha,\mu,\sigma>0$ and $B=(B(t),t\geq 0)$ is a scalar Brownian motion. The CIR model is mean-reverting and due to its square root diffusion factor, it can also avoid possible negative rates.
We observe in SDE (1) that the volatility term $\sigma$ is assumed constant. However, as supported by empirical studies, volatility is not constant but exhibits empirical features widely known as volatility skews and smiles that are prevalent in option markets (see, e.g., [7, 6]). To provide adequate description of the evolution of volatility skews and smiles, one of the main schools of thought proposed based on empirical findings is to model volatility as a time-lagged dependent variable. For instance, Arriojas et al. in [1] proposed the delay Black-Scholes model where the drift depends on a time-lagged dependent variable and, also, the volatility term is a function of a time-lagged dependent variable. This model is described by the dynamics
@@
on $t\in[0,T]$ with initial value $\xi(t)\in[-\tau,0]$, where $\xi:\Omega\rightarrow C([-\tau,0]:\mathbb{R})$, $\mu,a,b>0$, $\tau=\max(a,b)$, x(ta)x(t-a) and x(tb)x(t-b) denote delays in $x(t)$, $g:\mathbb{R}\rightarrow\mathbb{R}$ is a continuous function and $B=(B(t),t\geq 0)$ is a scalar Brownian motion. The authors showed that the delay Black-Scholes model maintains the no-arbitrage property and the completeness of the market with the correct volatility skews and smiles. Furthermore, Mao and Sabanis in [13] introduced the delay geometric Brownian motion described by the dynamics
@@
on $t\geq 0$ with the initial value $\xi(t)$ on $t\in[-\tau,0]$, where $\tau,r>0$, $\xi:\Omega\rightarrow C([-\tau,0]:\mathbb{R})$, $B=(B(t),t\geq 0)$ is a scalar Brownian motion, the volatility function $v$ depends on $x(t-\tau)$ and $x(t-\tau)$ denotes delay in $x(t)$. The authors studied the quantitative properties of this model where $v$ is merely local Lipschitz continuous and bounded. The authors provided numerical evidence to justify that the system of type (3) is a rich alternative model for an asset price process in a complete market characterised by volatility skews and smiles. The reader may also consult (e.g., [16, 9]) for financial models with features of past dependency.
In recent years, more empirical studies have shown that the most successful continuous-time models for interest rates are those models that allow the volatility of interest changes to be highly sensitive to the level of the rates (see, e.g., [3, 14]). This motivated Wu et al. in [15] to extend SDE (1) to include super-linear diffusion term described by the dynamics
@@
on $t\geq 0$, where $\theta>1$. One notable unique feature of SDE (4) is that the solution $x(t)$ is a highly sensitive mean-reverting process. The authors established the convergence in probability of the EM solutions to the true solution. They justified the convergence result within Monte Carlo simulations to value the expected payoff of a bond and a barrier option. To further capture volatility skews and smiles and high non-linearities in the rate, the authors in [4] extended the generalised Ait-Sahalia interest rate model to include a volatility as a function of a time-lagged dependent variable described by
@@
on $t\geq 0$ with the initial value $\xi(t)$ on $t\in[-\tau,0]$, where $\alpha_{-1},\alpha_{0},\alpha_{1},\alpha_{2},\tau>0$, $\xi:\Omega\rightarrow C([-\tau,0]:\mathbb{R})$, $\gamma,\theta>1$ and $B=(B(t),t\geq 0)$ is a scalar Brownian motion, $v$ is a function of $x(t-\tau)$ and $x(t-\tau)$ denotes delay in $x(t)$. Under a monotone condition and the assumption that $v$ is local Lipschitz continuous and bounded, the authors proved the strong convergence of the truncated EM solutions to the true solution of SDDE (5) and justified that the strong convergence result can be applied to value a bond and a barrier option.
Therefore, in other to account for high non-linearities in the rates as well as the evolution of volatility skews and smiles, we consider it necessary to reformulate SDE (4) as SDDE with super-linearly growing drift and diffusion coefficients described by the dynamics
@@
on $t\geq 0$ with the initial value $\xi(t)$ on $t\in[-\tau,0]$, where $\tau>0$, $\xi:\Omega\rightarrow C([-\tau,0]:\mathbb{R})$, $\theta,r>0$ and $\gamma>1$. We observe that both the drift factor $x^{\gamma}$ and the diffusion factor $x^{\theta}y^{r}$ of SDDE (6) are growing super-linearly and thus violate the global Lipschitz and linear growth conditions. This is further complicated by the presence of the unbounded time-lagged dependent variable $y$. Therefore, it can be very challenging to obtain the solution of SDDE (6) by an analytical closed-form formula. To the best of our knowledge, there exist no relevant literature for numerical treatment of SDDE (6) either in the strong sense or weak sense. In this case, we recognise the need to examine the feasibility of the system of SDDE (6) from a viewpoint of financial applications. Therefore, we need an efficient numerical method with fast computational performance to estimate the solution. However, in most real-world applications, the explicit EM method is preferred to the implicit type due to its simple algebraic structure, cheap computational cost and acceptable convergence rate. It is well-known in [8] that the explicit EM scheme diverges in the strong mean-square sense at finite point for SDEs with super-linearly growing coefficient terms. In this work, we aim to construct a variant of the truncated EM method developed in [11] to estimate the true solution of SDDE (6) and show that the truncated EM solutions converge to the true solution in probability when the step size is sufficiently small. The remainder of the paper is organised as follows: We explore mathematical notations in Section 2. In Section 3, we study properties of the true solution of SDDE (6). We construct the truncated EM techniques to approximate SDDE (6) and study properties of the truncated EM solutions in Section 4. In Section 5, we show that the truncated EM solutions converge to the true solution of SDDE (6) in probability. We also provide illustrative numerical examples to support the convergence result and justify the result via efficient use of the Monte Carlo method to value a bond and a lookback put option in this section .
@@ 2 Mathematical preliminaries
There exist constants $D>0$ and $\ell\in(0,1]$ such that for all $\tau\leq s\leq t\leq 0$, the initial function $\xi$ satisfies
@@
We introduce the following lemma for later use.
@@
Let Assumption 2.1 hold. For any $R>0$, there exists a constant $G_{R}>0$ such that the coefficients of the SDDE (6) satisfy
$|f(x)-f(\bar{x})|+|g(x,y)-g(\bar{x},\bar{y})|\leq G_{R}\big(|x-\bar{x}|+|y-\bar{y}|\big)$ (12)
for all $x,y,\bar{x},\bar{y}\in\mathbb{R}$ with $\max\{|x|,|\bar{x}|,|y|,|\bar{y}|\}\leq R$.
@@
We only prove the result for $x\leq\bar{x}$ and $y\leq\bar{y}$; the general case follows similarly. First observe that
$(f(x)-f(\bar{x}))+(g(x,y)-g(\bar{x},\bar{y}))$
$=\alpha(\mu-x^{\gamma})-\alpha(\mu-\bar{x}^{\gamma})+\sigma x^{\theta}y^{r}-\sigma\bar{x}^{\theta}\bar{y}^{r}$
$=-\alpha(x^{\gamma}-\bar{x}^{\gamma})+\sigma(x^{\theta}y^{r}-\bar{x}^{\theta}\bar{y}^{r}).$
Using Young inequality, we have
$x^{\theta}y^{r}\leq x^{2\theta}+y^{2r},\quad\bar{x}^{\theta}\bar{y}^{r}\leq\bar{x}^{2\theta}+\bar{y}^{2r}.$
Hence,
$x^{\theta}y^{r}-\bar{x}^{\theta}\bar{y}^{r}\leq(x^{2\theta}-\bar{x}^{2\theta})+(y^{2r}-\bar{y}^{2r}).$
It follows that
$(f(x)-f(\bar{x}))+(g(x,y)-g(\bar{x},\bar{y}))$
$\leq-\alpha(x^{\gamma}-\bar{x}^{\gamma})+\sigma(x^{2\theta}-\bar{x}^{2\theta})+\sigma(y^{2r}-\bar{y}^{2r}).$
Applying the mean value theorem to each term yields
$|x^{\gamma}-\bar{x}^{\gamma}|$ $\leq\gamma\big(|x|^{\gamma-1}+|\bar{x}|^{\gamma-1}\big)|x-\bar{x}|,$
$|x^{2\theta}-\bar{x}^{2\theta}|$ $\leq 2\theta\big(|x|^{2\theta-1}+|\bar{x}|^{2\theta-1}\big)|x-\bar{x}|,$
$|y^{2r}-\bar{y}^{2r}|$ $\leq 2r\big(|y|^{2r-1}+|\bar{y}|^{2r-1}\big)|y-\bar{y}|.$
Therefore,
$|f(x)-f(\bar{x})|+|g(x,y)-g(\bar{x},\bar{y})|$
$\leq\Big[\alpha\gamma(|x|^{\gamma-1}+|\bar{x}|^{\gamma-1})+2\theta\sigma(|x|^{2\theta-1}+|\bar{x}|^{2\theta-1})\Big]|x-\bar{x}|$
$\quad+2r\sigma(|y|^{2r-1}+|\bar{y}|^{2r-1})|y-\bar{y}|.$
Since $|x|,|\bar{x}|,|y|,|\bar{y}|\leq R$, all coefficients are bounded. Hence there exists a constant $G_{R}>0$ such that
$|f(x)-f(\bar{x})|+|g(x,y)-g(\bar{x},\bar{y})|\leq G_{R}(|x-\bar{x}|+|y-\bar{y}|).$
This completes the proof. ∎
@@ 3.1 Existence of positive solution
As $k\rightarrow\infty$, we have $\eta_{\infty}>T$ almost surely. Repeating this procedure for $t_{1}\in[0,\infty)$, it is easy to see that $\mathbb{P}(\eta_{\infty}\leq\infty)\rightarrow 0$ as $k\rightarrow\infty$. This implies that $\eta_{\infty}=\infty$ almost surely and hence we must have $\eta_{e}=\infty$ almost surely. Therefore, the solution remains strictly positive and finite almost surely on the interval $[0,T]$, that is,
$\mathbb{P}\big(0<x(t)<\infty\text{ for all }t\in[0,T]\big)=1.$
@@ 3.2 Boundedness
$\mathbb{P}(1/k<x(t)<k)\geq 1-\Big[\xi(0)^{\beta}-1-\beta\log(\xi(0))+\bar{K}_{0}\Big]\Big[\frac{1}{(1/k)^{\beta}-1+\beta\log(k)}+\frac{1}{k^{\beta}-1-\beta\log(k)}\Big],$
@@ 5 Convergence analysis
In this section, we study the finite-time convergence of the truncated EM solutions to the true solution of SDDE (6). Further, we show that the truncated EM solutions converge to the true solution of SDDE (6) in probability. We perform simulation examples to support the findings and justify the convergence result for valuing some option contracts.
@@ 5.2 Convergence in probability
$\frac{2V(\xi(0))+K_{0}T+K_{1}T}{V(k)\wedge V(1/k)}\leq\frac{\epsilon}{2}$ (39)
@@
$\frac{K_{3}D\Delta^{\ell}+c_{p}(K_{2}+K_{3})\Delta^{1/2}\psi(\Delta)T}{V(1/k)\wedge V(k)}+\frac{K_{8}(\Delta^{\ell}\vee\Delta^{p/2}(\psi(\Delta))^{p})}{\lambda}\leq\frac{\epsilon}{2}.$ (40)
@@
$\mathbb{P}\Big(\sup_{0\leq t\leq T}|x_{\Delta}(t)-x(t)|^{p}\geq\lambda\Big)\leq\epsilon,$ (41)
@@ Acknowledgement
The author would like to acknowledge the financial support from the Heilbronn Institute for Mathematical Research (HIMR) and the UKRI/EPSRC Additional Funding Programme for Mathematical Sciences. The author would also like to thank Prof Mao Xuerong for the discussion about this work.